How much heat is required to convert solid ice with a mass of 525 g and at a temperature of-23.5 °C to liquid water at a temperature of 48.5 °C? (The specific heat of ice is cice = 2100 /kgK, the specific heat of water is cwater-4186.8 J/kgK, and the heat of fusion for water is: Lf 334 kJ/kg.) Submit Answer Tries 0/12
Answer #1

Mass of solid ice = 525g or 0.525kg

Initial temperature = -23.5

Final temperature = 48.5

NOw,

To warm the ice to its melting point, the heat is

Q = ms \Delta T

Q = (2100)(0.525)(0-(-23.5))

Q = 25908.8J

Heat required to melt the ice is

(334KJ/kg)(0.525kg) = 175.35kJ or  175350J

Now the heat required to warm the melted ice to 48.5 degree is

(4186.8)(0.525)(48.5-0) = 106606J

Thus the total heat required is

25908.8 + 175350 + 106606 = 307865J

or

307.865 kJ